Logic007:

Form a suitable differential eqn using y=Ax^2+Bx^5

Logic007:

\(y'=2Ax+5Bx^4\) \(y''=2A+20Bx^3\)

Logic007:

can I divide \(\frac{y'}{y''}\) ?

nuts:

yolo if you still need help lmk i can do it

nuts:

and yeah you can, it depends on what kind of differential equation they want. if you divide i don't think it's linear.

Logic007:

They didnt specify what kind of differential eqn. Yes.I do need help XD

Logic007:

yep,I trying to figure out how to eliminate A and B

nuts:

this equation is the solution to a class of differential equations called the euler I believe

nuts:

if you're familiar with differential equations it should be enough for you to come up with the equation that you need, but if not i can probably look further

Logic007:

Not really familiar XD but thanks for the info Is there any other way besides using euler eqns

nuts:

depends on if you need to remove A and B. But honestly just reading the first few lines of that should be enouhg

Logic007:

The question ask to eliminate A and B. But do u know how to eliminate those constants?

nuts:

yes, just formulate the appropriate euler differential equation by working backwards from the link I gave you

Logic007:

aite,might take me a while to digest it.

nuts:

well there is an easier way I guess

nuts:

y=Ax^2+Bx^5 y'=2Ax+5Bx^4 y''=2A+20Bx^3 y'''=60Bx^2 y4=120Bx y5=120B (y5=y''''') y''=2A+20Bx^3 y''=2A+20(y'''''/120)x^3 A=y''/2-y'''''/12*x^3 y=(y''/2-y'''''/12*x^3)*x^2+y'''''/120*x^5 y=y''/2*x^2-y'''''/12*x^5+y'''''/120*x^5 y=y''/2*x^2-(9/120)*y'''''*x^5

nuts:

granted i don't think that was easier cause that took more thinking than solving the euler differential equation did

Logic007:

lol,yes XD I think I will just use euler eqns If I got any other solutions,I share it here ^

Logic007:

Thanks bro ^

nuts:

yup no problem

Logic007:

We cant differentiate y up to 5 times because there are only 2 arbitrary constants Since we were given 2 arbitrary constants in the solution , the original diff. eqn. will hv at most 2 derivative-that the most derivatives

Logic007:

\((1) \quad y = Ax^2 + B x^5 \\ \, \\ (2) \quad y' = 2A x + 5Bx^4 \\ \, \\ (3) \quad y'' = 2A + 20B x^3\) Multiply equation (2) by x, and equation(3) by x^2 \((1) \quad y = Ax^2 + B x^5 \\ \, \\ (2) \quad xy' = 2A x^2 + 5Bx^5 \\ \, \\ (3) \quad x^2 y'' = 2Ax^2 + 20B x^5\) Multiply equation (1) by 2 \((1) \quad 2y = 2Ax^2 + 2B x^5 \\ \, \\ (2) \quad xy' = 2A x^2 + 5Bx^5 \\ \, \\ (3) \quad x^2 y'' = 2Ax^2 + 20B x^5\) Subtract, equation (2) - (1) and equation (3) - (1) \((1) \quad 2y = 2Ax^2 + 2B x^5 \\ \, \\ (2) \quad xy' - 2y = 0 + 3B x^5 \\ \, \\ (3) \quad x^2 y'' - 2y = 0 + 18 Bx^5\) multiply equation(2) by 6 \((1) \quad 2y = 2Ax^2 + 2B x^5 \\ \, \\ (2) \quad 6xy' -12y = 0 + 18B x^5 \\ \, \\ (3) \quad x^2 y'' - 2y = 0 + 18 Bx^5\) Subtract equation(3) - equation(2) \((1) \quad 2y = 2Ax^2 + 2B x^5 \\ \, \\ (2) \quad 6xy' -12y = 0 + 18B x^5 \\ \, \\ (3) \quad x^2 y'' - 2y - 6xy' + 12y = 0 + 0\) Then equation (3) gives us \(x^2 y'' -6xy' + 10y = 0\)

nuts:

that's a good point that i didn't realize @Logic007 . Glad you figured it out lol

Logic007:

yeah,I just realized it XD I got another question related to separable method

nuts:

those are fun and easy... is this for an ordinary differential equations class?

Logic007:

yep,that's right I just finish sem 1 finals rn,it is sem break for me But I decided to to finish differential eqn b4 I enter sem 2 XD

nuts:

I had a lot of fun with differential equations, though our school teaches it as applied math so we don't go in detail into the theory

Logic007:

Owh I c. http://prntscr.com/m39l3d This are the chapters for diff eqn subject

Logic007:

These*

nuts:

ahh brings back memories from 3 years ago haha

Logic007:

haha XD